OpenStudy (anonymous):

Combinations: A schools academic consists of 7 students but only 3 can participate. How many different 3 person teams can be formed from 7 students? so 7! --- (7-3)!3! ??

jimthompson5910 (jim_thompson5910):

good, just evaluate that and you're done

OpenStudy (anonymous):

I got 105 tho.. because 7x6x5x4x3x2 ------------ 4x3x2x3x2 I canceled out the 2,3,4,2 ehh im confused /:

jimthompson5910 (jim_thompson5910):

it's not 105

jimthompson5910 (jim_thompson5910):

you are forgetting about the 3! in the denominator

OpenStudy (anonymous):

is it 35?

jimthompson5910 (jim_thompson5910):

7! --------- (7-3)!3! 7! --------- 4!3! 7*6*5*4! --------- 4!3! 7*6*5 --------- 3! 7*6*5 --------- 3*2*1 35 So you are correct, it's 35

jimthompson5910 (jim_thompson5910):

Faster way to do it There are 7 ways to fill slot 1 There are 6 ways to fill slot 2 There are 5 ways to fill slot 3 so there are 7*6*5 = 210 ways to do this if ordered mattered....but it does NOT matter So you have to divide by 3! = 3*2*1 = 6 to get 210/6 = 35

jimthompson5910 (jim_thompson5910):

This is why you'll see n C r = (n P r)/( r! )

OpenStudy (anonymous):

thank you!! :)

jimthompson5910 (jim_thompson5910):

sure thing